A box measures 3.0 ft x 2.0 ft x 1.50 ft. Each book measures 25.0 cm x 10.0 cm x 5.0 cm. What is the maximum number of this books that this box can contain?
1
Expert's answer
2020-09-30T12:25:36-0400

We will convert feet to centimeters by multiplying dimension in feet by k=30.48 cm/ft.

Volume of the box:


"V_\text{cm}=kAkBkC=k^3ABC=\\
=30.48^3\cdot3\cdot2\cdot1.5=254851.6\text{ cm}^3."

Volume of a book:


"v_\text{cm}=abc=\\
=25\cdot10\cdot5=1250\text{ cm}^3."

The integer part of the ratio "volume of the box" to "volume of one book" will be number of these books this box can contain is


"n=\frac{V}{v}=\frac{254851.6}{1250}=203.88"

Thus, the maximum number of books is 203.